OpenStudy (anonymous):

The initial profit is $147. The profit increases by 289% every 12 years. What is the yearly growth factor?

OpenStudy (anonymous):

yearly factor is:(3.89)^1/12 = 1.11986 \[3.89^{1/12} = \sqrt[12]{3.89} \approx 1.11986\] yearly growth rate is 1.11986 * 100% - 100% = 11.986%

OpenStudy (anonymous):

how do you get the 3.89?

OpenStudy (anonymous):

oh, sorry.... use 2.89 instead of 3.89 please.

OpenStudy (anonymous):

the 3.89 was right. I didn't know how you got it

OpenStudy (anonymous):

the initial profit is 100%, since it's increase which means after 12 years it increased 289% plus the original 100% you get 389%, and 389% = 389/100 = 3.89